Suppose at a price of $50, Yoshi’s Jazz Bar sells 20 tickets to its nightly jazz performance and at a price
of $40, it sells 25 tickets. Based on this information, answer the following using the ARC method:
A) What is the price elasticity of demand for tickets to Yoshi’s Jazz Bar?
B) What is the degree of elasticity for the tickets?
C) What would happen to sales if Yoshi’s Jazz Bar lowered its ticket price by 2% to its sales (numeric
answer)?
D) What happens to the total revenue when they drop their price from $50 to $40?
1
Expert's answer
2017-03-06T15:14:06-0500
A)
Q0=50
P0=20
Q1=40
P1=25
ARC elasticity of demand = (∆Q/∆P)*(P1+P0)/(Q1+Q0)
ARC elasticity of demand = -10/5 *45/90 = -1
B) Unitary elastic demand
C) Sales will increase by 2%.
For P=25, Q=40: if ∆P= -0.5, than ∆Q=0.8
D) The total revenue won’t change
TR0=50*20=1000
TR1=40*25=1000
∆TR=0